What is the mass (in g) of a sample of SF4 that contains 5.5E+23 fluorine atoms?
1
Expert's answer
2020-09-14T08:25:17-0400

Solution:

1) Let's calculate the number of moles of fluorine:

Moles of fluorine = n(F) = 5.5×1023 atoms × (1 mol / 6.022×1023 atoms) = 0.9133 mol


2) Let's calculate the number of moles of SF4:

Schematic equation: SF4 --> 4F

n(SF4) = n(F) / 4 (according to the equation)

n(SF4) = 0.9133 mol / 4 = 0.2283 mol


3) Let's calculate the mass of a sample of SF4:

Moles of SF4 = Mass of SF4 / Molar mass of SF4

The molar mass of SF4 is 108.07 g/mol.

Thus:

Mass of SF4 = Moles of SF4 × Molar mass of SF4 = (0.2283 mol) × (108.07 g/mol) = 24.67 g

The mass of a sample of SF4 is 24.67 g.


Answer: The mass of a sample of SF4 is 24.67 g.